# Let the Expert Stick to His Last: Expert-Specialized Fine-Tuning for Sparse Architectural Large Language Models

### v1

@@ Abstract
Parameter-efficient fine-tuning (PEFT) is crucial for customizing Large Language Models (LLMs) with constrained resources. Although there have been various PEFT methods for dense-architecture LLMs, PEFT for sparse-architecture LLMs is still underexplored. In this work, we study the PEFT method for LLMs with the Mixture-of-Experts (MoE) architecture and the contents of this work are mainly threefold: (1) We investigate the dispersion degree of the activated experts in customized tasks, and found that the routing distribution for a specific task tends to be highly concentrated, while the distribution of activated experts varies significantly across different tasks. (2) We propose Expert-Specialized Fine-Tuning, or ESFT, which tunes the experts most relevant to downstream tasks while freezing the other experts and modules; experimental results demonstrate that our method not only improves the tuning efficiency, but also matches or even surpasses the performance of full-parameter fine-tuning. (3) We further analyze the impact of the MoE architecture on expert-specialized fine-tuning. We find that MoE models with finer-grained experts are more advantageous in selecting the combination of experts that are most relevant to downstream tasks, thereby enhancing both the training efficiency and effectiveness.
@@ 2.1 Parameter-efficient fine-tuning for dense architectural LLMs
The goal of parameter-efficient fine-tuning (Han et al., 2024) is to efficiently customize LLMs for downstream tasks, while existing studies primarily focus on dense architectural LLMs. PEFT methods for dense models can generally be categorized into three approaches: (1) Adding new parameters: methods of this kind fix the existing model parameters and fine-tune the model on a small number of newly added parameters. Adapter (Houlsby et al., 2019; Pfeiffer et al., 2020; He et al., 2021; Wang et al., 2022) and Soft Prompt (Li and Liang, 2021; Liu et al., 2021; Zhang et al., 2023b; Lester et al., 2021) are two typical representatives of this category of methods. (2) Selecting existing parameters: methods of this type fine-tune a limited part of existing parameters, while keeping the majority of the other parameters fixed. Based on whether the trainable parameter space is continuous, these methods can generally be divided into structured training (Guo et al., 2020; Gheini et al., 2021; He et al., 2023; Vucetic et al., 2022) and unstructured training (Liao et al., 2023; Ansell et al., 2021; Sung et al., 2021; Xu et al., 2021). (3) Applying low-rank adaptation: LoRA (Hu et al., 2021; Fomenko et al., 2024) is a widely-used PEFT method, which decomposes the origin weight matrices into low-rank components. Subsequent works (Zhang et al., 2023a; Ding et al., 2023; Lin et al., 2024; Liu et al., 2023) have introduced numerous improvements to the original LoRA method. However, the study of PEFT in sparse models is still scarce. In this work, we select and tune part of the experts based on their downstream task affinity, which is a unique selection dimension exclusive to the sparse MoE architecture.
@@ 4.1.2 Tasks for Model Adaptation
To measure model performance, for the text-to-JSON task, we calculate the exact match between model output and reference answer; for other tasks, we employ GPT-4 to score model output between 0 and 10 given reference answer444The exact version we use is gpt-4-1106-preview. The evaluation instructions are in Appendix G..
@@ 4.3 Backbone Model and Training Settings
For hyperparameter settings, all methods use a batch size of 32 and a sequence length of 4096 for training. For every task, we set the maximum steps of training to 500, and evaluate the model every 100 steps. The learning rates are set to 3e-5, 1e-4, and 1e-5 for FFT, LoRA, and ESFT, respectively, based on a hyperparameter search in {1e-5, 3e-5, 1e-4, 3e-4}. The LoRA rank is set to 8 and scaling is set to 2, following Hu et al. (2021). The threshold is set to 0.1 for ESFT-Gate and 0.2 for ESFT-Token, respectively. §6.2 shows how we determine the threshold for ESFT666we plan to release the ESFT training and evaluation script, the vanilla backbone model, and all trained checkpoints on different tasks..
@@ 5.1 Benchmark Performance Results
For general ability evaluation, ESFT consistently outperforms FFT and LoRA by showing less performance degradation. As illustrated in Table 2, ESFT-token performs better than ESFT-gate, with average scores of 61.5 and 60.6, respectively. The results demonstrate a wide range of retention in tasks such as TriviaQA and IFEval, surpassing FFT’s 58.8 and LoRA’s 59.1. Both methods retain performance better than LoRA and FFT, highlighting their effectiveness in maintaining general task performance777 We further investigate Math and Code performance of the models trained on specialized tasks in Appendix H. FFT and LoRA exhibit even more severe degradation, while ESFT shows a minimal performance drop. . Analyses in §6.3 indicate that such degradation on general tasks for FFT and LoRA may result from training shared parameters.
@@ 6.4 Analysis of Key Modules in ESFT
We use the fine-grained segmented DeepSeek-V2 MoE model as our backbone. To prove the effectiveness of such fine-grained expert segmentation, we use greedy search (as detailed in Appendix B) to group experts, simulating coarse-grained expert segmentation. Experts in the same group share the same gate for each token, initialized by the average of the original gates’ vector. We conduct experiments in the Math domain as an example. Results in Figure 7 show that as the group size increases, the performance of our method decreases more severely than FFT, while the training cost (i.e., the average number of experts used) becomes larger. These findings show that our method, as well as even more effective LLM customization, highly relies on a fine-grained segmented LLM architecture that has more specialized experts.
@@ Appendix H Evaluating Math and Code as General Tasks
We further investigate the Math and Code performance of models trained on adaptation tasks (i.e., Intent, Summary, Law, Translation), as these domains reflect the model’s general ability if not specifically trained on them. We report numbers with the setting of training on only downstream task data. Results in Table 8 show that FFT and LoRA would lead to significant performance drops in the Math and Code domain, having average performance drops of 9.0 and 12.4, respectively. Notably, our ESFT method retains performance significantly better compared to FFT and LoRA, with an average performance drop of less than 1.

### v2

@@ Abstract
Parameter-efficient fine-tuning (PEFT) is crucial for customizing Large Language Models (LLMs) with constrained resources. Although there have been various PEFT methods for dense-architecture LLMs, PEFT for sparse-architecture LLMs is still underexplored. In this work, we study the PEFT method for LLMs with the Mixture-of-Experts (MoE) architecture and the contents of this work are mainly threefold: (1) We investigate the dispersion degree of the activated experts in customized tasks, and found that the routing distribution for a specific task tends to be highly concentrated, while the distribution of activated experts varies significantly across different tasks. (2) We propose Expert-Specialized Fine-Tuning, or ESFT, which tunes the experts most relevant to downstream tasks while freezing the other experts and modules; experimental results demonstrate that our method not only improves the tuning efficiency, but also matches or even surpasses the performance of full-parameter fine-tuning. (3) We further analyze the impact of the MoE architecture on expert-specialized fine-tuning. We find that MoE models with finer-grained experts are more advantageous in selecting the combination of experts that are most relevant to downstream tasks, thereby enhancing both the training efficiency and effectiveness. Our code is available at https://github.com/deepseek-ai/ESFT.
@@ 2.1 Parameter-efficient fine-tuning for dense architectural LLMs
The goal of parameter-efficient fine-tuning (Han et al., 2024) is to efficiently customize LLMs for downstream tasks, while existing studies primarily focus on dense architectural LLMs. PEFT methods for dense models can generally be categorized into three approaches: (1) Adding new parameters: methods of this kind fix the existing model parameters and fine-tune the model on a small number of newly added parameters. Adapter (Houlsby et al., 2019; Pfeiffer et al., 2020; He et al., 2021; Wang et al., 2022) and Soft Prompt (Li and Liang, 2021; Liu et al., 2021; Zhang et al., 2023b; Lester et al., 2021) are two typical representatives of this category of methods. (2) Selecting existing parameters: methods of this type fine-tune a limited part of existing parameters, while keeping the majority of the other parameters fixed. Based on whether the trainable parameter space is continuous, these methods can generally be divided into structured training (Guo et al., 2020; Gheini et al., 2021; He et al., 2023; Vucetic et al., 2022) and unstructured training (Liao et al., 2023; Ansell et al., 2021; Sung et al., 2021; Xu et al., 2021). (3) Applying low-rank adaptation: LoRA (Hu et al., 2021; Fomenko et al., 2024) is a widely-used PEFT method, which decomposes the origin weight matrices into low-rank components. Subsequent works (Zhang et al., 2023a; Ding et al., 2023; Lin et al., 2024; Liu et al., 2023) have introduced numerous improvements to the original LoRA method. However, the study of PEFT in sparse models is still scarce. In this work, we select and tune part of the experts based on their downstream task affinity, as a unique selection dimension exclusive to the sparse MoE architecture.
@@ 4.1.2 Tasks for Model Adaptation
To measure model performance, for the text-to-JSON task, we calculate the exact match between model output and reference answer; for other tasks, we employ GPT-4 to score model output between 0 and 10 given reference answer444The exact version we use is gpt-4-1106-preview. The evaluation instructions are in Appendix G.. All evaluations use few-shot examples.
@@ 4.3 Backbone Model and Training Settings
For hyperparameter settings, all methods use a batch size of 32 and a sequence length of 4096 for training. For every task, we set the maximum steps of training to 500, and evaluate the model every 100 steps. The learning rates are set to 3e-5, 1e-4, and 1e-5 for FFT, LoRA, and ESFT, respectively, based on a hyperparameter search in {1e-5, 3e-5, 1e-4, 3e-4}. The LoRA rank is set to 8 and scaling is set to 2, following Hu et al. (2021). The threshold is set to 0.1 for ESFT-Gate and 0.2 for ESFT-Token, respectively. §6.2 shows how we determine the threshold for ESFT.
@@ 5.1 Benchmark Performance Results
For general ability evaluation, ESFT consistently outperforms FFT and LoRA by showing less performance degradation. As illustrated in Table 2, ESFT-token performs better than ESFT-gate, with average scores of 61.5 and 60.6, respectively. The results demonstrate a wide range of retention in tasks such as TriviaQA and IFEval, surpassing FFT’s 58.8 and LoRA’s 59.1. Both methods retain performance better than LoRA and FFT, highlighting their effectiveness in maintaining general task performance666 We further investigate Math and Code performance of the models trained on specialized tasks in Appendix H. FFT and LoRA exhibit even more severe degradation, while ESFT shows a minimal performance drop. . Analyses in §6.3 indicate that such degradation on general tasks for FFT and LoRA may result from training shared parameters.
@@ 6.4 Analysis of Key Modules in ESFT
We use the fine-grained segmented DeepSeek-V2 model as our backbone. To demonstrate t the effectiveness of this fine-grained segmentation, we use greedy search (as detailed in Appendix B) to group experts, simulating coarse-grained segmentation. Experts in the same group share the average affinity score. We maintain the computational cost by selecting a constant 1/8 of experts for each token. Experiment results of the Math domain in Figure 7 show that as the group size increases, our method’s performance decreases more severely than FFT, while the training cost (i.e., trainable experts) rises. These findings indicate that our method, and even effective LLM customization, highly rely on a fine-grained segmented LLM architecture with more specialized experts.
@@ Appendix H Evaluating Math and Code as General Tasks
We investigate the Math and Code performance of models trained on adaptation tasks (i.e., Intent, Summary, Law, Translation), as these domains reflect the model’s general ability if not specifically trained on them. We report numbers with the setting of training on only downstream task data. Results in Table 8 show that FFT and LoRA would lead to significant performance drops in the Math and Code domain, having average performance drops of 9.0 and 12.4, respectively. Notably, our ESFT method retains performance significantly better compared to FFT and LoRA, with an average performance drop of less than 1.